Joshua drives a jet ski on a sunny day 100 km away. The current speed of the water is 50 mph. He drives the jet ski against the current for one hour. Then he makes the same trip back with the current, but this time it takes him 40 minutes. How fast is he traveling in still water?
Answer by stanbon(75887)

Joshua drives a jet ski on a sunny day 100 km away. The current speed of the water is 50 mph. He drives the jet ski against the current for one hour. Then he makes the same trip back with the current, but this time it takes him 40 minutes. How fast is he traveling in still water?
---
Against current DATA:
distance = 100km ; time = 1 hr; rate = 100 km/hr.
----------------------
With current DATA:
distance = 100 km ; time = (2/3)hr ; rate = 100/(2/3) = 150 km/hr.
-----------------------
Equation:
r - 50 = 100 km/h
r + 50 = 150 km/h
-----------------------
Add and solve for "r":
2r = 250
r = 125 km/h (his speed in still water)
==========================================
Cheers,
Stan H.
===========